A poll was taken of 100 students at a commuter campus to find out how they got to campus. The results were as follows:
32 said they drove alone.
27 rode in a carpool.
28 rode public transportation.
6 used both carpools and public transportation.
7 used both a carpool and sometimes their own cars.
8 used public transit as well as their own cars.
4 used all three methods.
How many used none of the above-mentioned means of transportation?

Answer by richwmiller(17219)

32 said they drove alone.
27 rode in a carpool.
28 rode public transportation.
6 used both carpools and public transportation.
7 used both a carpool and sometimes their own cars.
8 used public transportation as well as their own cars.
4 used all three methods.
Venn Diagram with known info

alone
7 carpools and alone
8 public transportation and alone
4 all three ways
total 32-19=13 =a only alone
carpool
7 carpools and alone
4 all three ways
6 carpool and bus
27-17=10 =b only carpool
public transportation
8 public transportation and alone
4 all three ways
6 carpool and public transportation
28-18=10 =c only public transportation
Shared 8+4+6+7=25
13+10+10=33
25+33=58
100-58=42 =none
Venn Diagram with calculated info

none of above might be walking, bike, skateboard, roller skates, or motorbike